OpenStudy (anonymous):

Which of the following statements is false? A. Whole numbers are never irrational. B. There are no integers that are whole numbers. C. There is only one whole number that is not a natural number. D. Natural numbers cannot be negative.

OpenStudy (anonymous):

B. A is true as whole numbers indeed are always rational -- e.g. \[\frac{0}{1}, \frac{1}{1}, \frac{2}{1}, ...\] B is true as the set of non-negative integers is actually identical to the whole numbers. C is true depending on the convention chosen for what defines a natural number. For natural numbers defined as follows... \[ 1, 2, 3, ... \] i.e. not including zero, then indeed it's difference from the set of whole numbers is \[\{0\} \] D is true because regardless of convention it is agreed upon that natural numbers are at least non-negative.